# ID-Blau: Image Deblurring by Implicit Diffusion-based reBLurring AUgmentation

### v1

@@ Abstract
Image deblurring aims to remove undesired blurs from an image captured in a dynamic scene. Much research has been dedicated to improving deblurring performance through model architectural designs. However, there is little work on data augmentation for image deblurring. Since continuous motion causes blurred artifacts during image exposure, we aspire to develop a groundbreaking blur augmentation method to generate diverse blurred images by simulating motion trajectories in a continuous space. This paper proposes Implicit Diffusion-based reBLurring AUgmentation (ID-Blau), utilizing a sharp image paired with a controllable blur condition map to produce a corresponding blurred image. We parameterize the blur patterns of a blurred image with their orientations and magnitudes as a pixel-wise blur condition map to simulate motion trajectories and implicitly represent them in a continuous space. By sampling diverse blur conditions, ID-Blau can generate various blurred images unseen in the training set. Experimental results demonstrate that ID-Blau can produce realistic blurred images for training and thus significantly improve performance for state-of-the-art deblurring models.
@@ 3 Proposed Method
This section presents the proposed ID-Blau that can turn a sharp image into a blurred version based on a pixel-wise blur condition map. We characterize blurs with their orientations and magnitudes in a continuous space, where we can sample blur conditions for reblurring a sharp image. With the modeling, we can simulate blur trajectories caused by continuous motion during exposure. It allows us to implicitly manipulate blur conditions to generate diverse blurred images unseen in the training set for data augmentation. In the data augmentation scenario, we can use ID-Blau to produce additional training data before optimizing deblurring models, which means that we do not need to run ID-Blau when optimizing deblurring models. The following details the computation of blur conditions and the training and sampling processes of ID-Blau for blurred image generation.
@@
3.1 Blur Conditions
During exposure, continuous motion from the camera or scene objects may cause blur artifacts to captured images. Thus, global and/or local blur trajectories can be found in such blurred images. To synthesize blurred images from their sharp counterparts, Nah et al. [18] proposed the GoPro dataset, approximating the continuous exposure through a blur accumulation function as
[…]
where $B$, $V(t)$, $T$, $N$, and $g$ denote the generated blurred frame, sharp frame at time $t$ collected by a high-speed camera, exposure time, number of sampled sharp frames, and camera response function, respectively. The accumulation function aggregates a sharp image sequence $\textbf{V}=\{V\left[1\right],...,V\left[N\right]\}$ to generate one blurred image $B$. The center sharp frame $V[\frac{N+1}{2}]$ is assigned to the ground-truth image $S$ for $B$.
where $f_{\theta}$ denotes the pre-trained flow estimation network [29], $\mathcal{F}=[u;v]\in\mathbb{R}^{H\times W\times 2}$ denotes the overall motion trajectories during exposure, where $H$ and $W$ are the image height and width, respectively. The tensor $\mathcal{F}$ records the horizontal and vertical motion trajectories in $u$ and $v\in\mathbb{R}^{H\times W}$.
@@ 3.1 Blur Conditions
For example, we show the distribution of the blur magnitude and orientation of GoPro in Figure 2, indicating that the magnitude histogram peaks towards the left side, meaning most images are not heavily blurred. In contrast, the orientation has slight peaks around 0, 180, and 360 degrees, implying more blurs tend to be horizontal, but cases in other orientations can also be found. Training our reblurring model on such diverse data can correlate a blurred image with its computed blur information implicitly in a continuous blur condition field and empower the model to generate blurred images unseen in the original data by sampling different blur magnitudes and orientations. Next, we describe optimizing the proposed diffusion-based reblurring model in ID-Blau.
@@ 3.2 ID-Blau
ID-Blau is based on a conditional diffusion model that takes a blur condition map $C$ and a sharp image $S$ as inputs to generate a blurred image $B$, as shown in Figure 3. Therefore, we can create various blur conditions to augment data with realistic reblurred images unseen in the training set. To expedite the inference, we adopt the denoising diffusion implicit model (DDIM) [26], utilizing the same training procedure as DDPM [12] but accelerating the inference via a non-Markovian inference process.
@@
Figure 5: Visuals of blurred images with local motions. ID-Blau can generate blurred images of moving objects, such as pedestrians.
@@
where $B_{0}=B$ and $\alpha_{t}\in(0,1)$ is a set of hyperparameters that controls the noise variance over a sequence of steps. Instead of iteratively adding Gaussian noise to $B$, the forward diffusion process can be reparameterized as
@@
$\begin{split}B_{t-1}=&\sqrt{\alpha_{t-1}}(\frac{B_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(B_{t},S,C,t)}{\sqrt{\alpha_{t}}})\\ &+\sqrt{1-\alpha_{t-1}-\sigma^{2}_{t}}\cdot\epsilon_{\theta}(B_{t},S,C,t)+\sigma_{t}z,\end{split}$ (7)
@@
Figure 4 shows examples of generating different blurred images from a sharp image $S$ by ID-Blau. That is, given a blur condition map $C=[x;y;z]$ and a sharp image $S$, ID-Blau can produce a blurred image $B=\textbf{ID-Blau}(S,C)$. Additionally, we create four blur condition maps based on $C$ for further demonstration, including unit horizontal or vertical blur orientations, $C^{1}=[1;0;z]$ and $C^{2}=[0;1;z]$, horizontally inverse $C$ as $C^{3}=[-x;y;z]$, and $C^{4}=[-x;y;2z]$ with twice the magnitude for $C^{3}$. Furthermore, ID-Blau can also simulate motion blurs caused by object movements, as shown in Figure 5. Besides, we have more examples to demonstrate the generalizability of ID-Blau by utilizing sharp images not in the training set to yield different blurred images in the supplementary material. These visuals attest that ID-Blau can effectively generate controllable and realistic blurred images. Thus, it can augment training data to improve image deblurring performance.
@@ 4.1 Implementation Details
we adopt the GoPro training set [18], which contains $2,103$ blurred and sharp images for training ID-Blau. The training also requires the blur condition maps computed based on the blurred images. We use a batch size of $32$ and randomly crop images with the size of $128\times 128$ for training $5,000$ epochs. We adopt the Adam optimizer and maintain a fixed learning rate of $1e^{-4}$. After optimizing ID-Blau, we utilize sharp images in the GoPro training set and create extra blur condition maps to produce additional $10,000$ $1280\times 720$ blurred images by randomly modifying blur orientations and magnitudes based on the originally computed blur condition maps. The newly generated $10,000$ blurred and sharp pairs can largely enrich the GoPro training set to enhance the performance of a deblurring model. ID-Blau has 9.5 million parameters, and it takes 3.4 seconds to generate a $1280\times 720$ blurred image on an Nvidia 3090.
@@ 4.2 Experimental Results
Figure 6 compares the visual qualities of the results with the four deblurring baselines (denoted “Baseline”) and their ID-Blau-powered versions (denoted “+ID-Blau”) on GoPro and HIDE. The results show that ID-Blau significantly enhances deblurring visual qualities compared to the Baselines. Besides, we demonstrate visuals of deblurring models on RealBlur-J in Figure 7. We also demonstrate deblurring results on real-world blurry images [42] in the supplementary material.
@@ 4.3 Ablation Studies
Table 2 shows the effect of augmenting blur orientations with different angles on deblurring performance. To this end, with the original blur magnitudes, we rotate the blur orientations $[x;y]$ by four fixed angles ($30^{\circ}$, $60^{\circ}$, $90^{\circ}$, and $120^{\circ}$) plus a randomly selected angle for each produced blurred image (denoted “Random”). The results indicate that blur augmentation with all specific/random angles achieves similar performance improvements over Baseline. This suggests no need to employ specific angles in blur condition maps for augmentation. For implementation convenience, we randomly alter blur orientations $[x;y]$ to $[-x;y]$, $[x;-y]$, or $[-x;-y]$ to augment blurred images with different orientations, achieving a similar performance improvement as well.
@@
In ID-Blau, we adopt the diffusion process in [12] to generate blurred versions of sharp images with specified blur conditions. However, without the diffusion process, we can still utilize the same reblurring model to learn the mapping from a sharp image plus a computed blur condition map to the corresponding blurred image with the typical L1 loss function. To analyze the impact of ID-Blau with or without the diffusion process on deblurring performance, in Table 3, we compare Baseline (the same as in Figure 9), “w/ ${}^{\dagger}$,” (ID-Blau without diffusion), and “w/ ID-Blau” (our model). The results show that “w/ ${}^{\dagger}$” and “w/ ID-Blau” improve the Baseline by $0.57$dB and $0.80$dB, respectively, and the diffusion model achieves $0.23$dB additional performance gain.
@@
Table 3: Effectiveness of using a diffusion model in ID-Blau on MIMO-UNet’s performance in PSNR (dB), where “${}^{\dagger}$” denotes the proposed ID-Blau without its diffusion process.
Model Baseline w/ ${}^{\dagger}$ w/ ID-Blau
@@
Table 4: Comparison between ID-Blau and BGAN regarding the DBGAN’s performance in PSNR (dB), where ${}^{\dagger}$ denotes the proposed ID-Blau without its diffusion process.
Model Baseline w/ BGAN w/ ${}^{\dagger}$ w/ ID-Blau
PSNR 30.43 31.10 31.57 31.67
@@
Besides, we compare the performance gains of ID-Blau and that of BGAN [42], a representative reblurring GAN for augmenting blurred images to improve the DBGAN deblurring model. To fairly compare ID-Blau with BGAN, we utilize ID-Blau to improve DBGAN’s deblurring performance. In Table 4, “Baseline” denotes DBGAN’s performance trained on GoPro, and “w/ BGAN” denotes DBGAN’s performance with BGAN-augmented data. “w/ ${}^{\dagger}$” and “w/ ID-Blau” denote DBGAN’s performance using ID-Blau-augmented data without and with the diffusion process, respectively. Note that GAN-based methods tend to generate unpredictable and uncontrollable blurred images, whereas ID-Blau can generate controllable results based on specified blur conditions. As a result, “w/ ${}^{\dagger}$” and “w/ ID-Blau” improve the PSNR performance of DBGAN by $1.14$dB and $1.24$dB. Thanks to the proposed continuous blur condition map with ID-Blau, the simplified “w/ ${}^{\dagger}$” without the diffusion process can still offer an additional gain of $0.47$dB over that with “w/ BGAN.” This verifies the effectiveness of ID-Blau compared to the GAN-based reblurring model BGAN.

### v2

@@ Abstract
Image deblurring aims to remove undesired blurs from an image captured in a dynamic scene. Much research has been dedicated to improving deblurring performance through model architectural designs. However, there is little work on data augmentation for image deblurring. Since continuous motion causes blurred artifacts during image exposure, we aspire to develop a groundbreaking blur augmentation method to generate diverse blurred images by simulating motion trajectories in a continuous space. This paper proposes Implicit Diffusion-based reBLurring AUgmentation (ID-Blau), utilizing a sharp image paired with a controllable blur condition map to produce a corresponding blurred image. We parameterize the blur patterns of a blurred image with their orientations and magnitudes as a pixel-wise blur condition map to simulate motion trajectories and implicitly represent them in a continuous space. By sampling diverse blur conditions, ID-Blau can generate various blurred images unseen in the training set. Experimental results demonstrate that ID-Blau can produce realistic blurred images for training and thus significantly improve performance for state-of-the-art deblurring models. The source code is available at https://github.com/plusgood-steven/ID-Blau.
footnotetext: * equal contribution
@@ 3 Proposed Method
This section presents the proposed ID-Blau, which turns a sharp image into a blurred version based on a pixel-wise blur condition map. We characterize blurs with their orientations and magnitudes in a continuous space, where we can sample blur conditions for reblurring a sharp image. With the modeling, we can simulate blur trajectories yielded by continuous motion during exposure. It allows us to implicitly manipulate blur conditions to generate diverse blurred images unseen during training for data augmentation. In the data augmentation scenario, we can use ID-Blau to produce additional training data before optimizing deblurring models, so we do not need to run ID-Blau when optimizing deblurring models. The following details the computation of blur conditions and the training and sampling processes of ID-Blau for blurred image generation.
3.1 Blur Conditions
During exposure, continuous motion from a camera or scene objects may cause blur artifacts to captured images. Thus, global and/or local blur trajectories can be found in such blurred images. To synthesize blurred images from their sharp counterparts, Nah et al. [18] proposed the GoPro dataset, approximating the continuous exposure through a blur accumulation function as
[…]
where $B$, $V(t)$, $T$, $N$, and $g$ denote the generated blurred frame, sharp frame at time $t$ collected by a high-speed camera, exposure time, number of sampled sharp frames, and camera response function, respectively. The accumulation function aggregates a sharp image sequence $\textbf{V}=\{V\left[1\right],...,V\left[N\right]\}$ to generate one blurred image $B$. The center sharp frame $V[\frac{N+1}{2}]$ is assigned to the ground-truth image $S$ for $B$, and $N$ is typically an odd number.
where $f_{\theta}$ denotes the pre-trained optical flow estimation network [29], $\mathcal{F}=[u;v]\in\mathbb{R}^{H\times W\times 2}$ denotes the overall motion trajectories during exposure, where $H$ and $W$ are the image height and width, respectively. The tensor $\mathcal{F}$ records the horizontal and vertical motion trajectories in $u$ and $v\in\mathbb{R}^{H\times W}$.
@@ 3.1 Blur Conditions
For example, we show the distribution of the blur magnitude and orientation of GoPro in Figure 2, indicating that the magnitude histogram peaks towards the left side, meaning most images are not heavily blurred. In contrast, the orientation has slight peaks around 0, 180, and 360 degrees, implying more blurs tend to be horizontal, but cases in other orientations can also be found. Training our reblurring model on such diverse data can correlate a blurred image with its computed blur information implicitly in a continuous blur condition field and empower the model to generate blurred images unseen in the original data by sampling different blur magnitudes and orientations. In the following, we describe optimizing the proposed diffusion-based reblurring model in ID-Blau.
@@ 3.2 ID-Blau
ID-Blau is developed based on a conditional diffusion model that takes a blur condition map $C$ and a sharp image $S$ as input to generate a blurred image $B$, as shown in Figure 3. Therefore, we can create various blur conditions to augment data with realistic reblurred images unseen in the training set. To expedite the inference, we adopt the denoising diffusion implicit model (DDIM) [26], utilizing the same training procedure as DDPM [12] but accelerating the inference via a non-Markovian inference process.
@@
where $B_{0}=B$ and $\{\alpha_{t}\in(0,1)\}$ is a set of hyperparameters that controls the noise variance over a sequence of steps. Instead of iteratively adding Gaussian noise to $B$, the forward diffusion process can be reparameterized as
@@
$\begin{split}B_{t-1}=&\sqrt{\overline{\alpha}_{t-1}}(\frac{B_{t}-\sqrt{1-\overline{\alpha}_{t}}\epsilon_{\theta}(B_{t},S,C,t)}{\sqrt{\overline{\alpha}_{t}}})\\ &+\sqrt{1-\overline{\alpha}_{t-1}-\sigma^{2}_{t}}\cdot\epsilon_{\theta}(B_{t},S,C,t)+\sigma_{t}z,\end{split}$ (7)
@@
Figure 4 shows examples of generating different blurred images from a sharp image $S$ by ID-Blau. That is, given a blur condition map $C=[x;y;z]$ and a sharp image $S$, ID-Blau can produce a blurred image $B=\textbf{ID-Blau}(S,C)$. Additionally, we create four blur condition maps based on $C$ for further demonstration, including unit horizontal or vertical blur orientations, $C^{1}=[1;0;z]$ and $C^{2}=[0;1;z]$, horizontally inverse $C$ as $C^{3}=[-x;y;z]$, and $C^{4}=[-x;y;2z]$ with twice the magnitude for $C^{3}$.
Besides, we have more examples to demonstrate the generalizability of ID-Blau by utilizing sharp images not in the training set to yield different blurred images in the supplementary material. These visuals attest that ID-Blau can effectively generate controllable and realistic blurred images. Thus, it can augment training data to improve image deblurring performance.
4 Experiments
footnotetext: The authors from the universities in Taiwan completed the experiments on the datasets.
@@ 4.1 Implementation Details
We adopt the GoPro training set [18], which contains $2,103$ blurred and sharp images for training ID-Blau. The training also requires the blur condition maps computed based on the blurred images. We use a batch size of $32$ and randomly crop images with the size of $128\times 128$ for training $5,000$ epochs. We adopt the Adam optimizer and maintain a fixed learning rate of $1e^{-4}$. After optimizing ID-Blau, we utilize sharp images in the GoPro training set and create extra blur condition maps to produce additional $10,000$ $1280\times 720$ blurred images by randomly modifying blur orientations and magnitudes based on the originally computed blur condition maps. The newly generated $10,000$ blurred and sharp pairs can largely enrich the GoPro training set to enhance the performance of a deblurring model. ID-Blau has 9.5 million parameters, and it takes 3.4 seconds to generate a $1280\times 720$ blurred image on an Nvidia 3090 graphics card.
@@ 4.2 Experimental Results
Figure 5 compares the visual qualities of the results with the four deblurring baselines (denoted by “Baseline”) and their ID-Blau-powered versions (denoted by “+ID-Blau”) on GoPro and HIDE. The results show that ID-Blau significantly enhances deblurring visual qualities compared to the Baselines. Besides, we demonstrate visuals of deblurring models on RealBlur-J in Figure 6. We also demonstrate deblurring results on real-world blurry images [42] in the supplementary material.
@@ 4.3 Ablation Studies
Table 2 shows the effect of augmenting blur orientations with different angles on deblurring performance. To this end, with the original blur magnitudes, we rotate the blur orientations $[x;y]$ by four fixed angles ($30^{\circ}$, $60^{\circ}$, $90^{\circ}$, and $120^{\circ}$) plus a randomly selected angle for each produced blurred image (denoted “Random"). The results indicate that blur augmentation with all specific/random angles achieves performance improvements similar to those of Baseline. This suggests no need to employ specific angles in blur condition maps for augmentation. For implementation convenience, we randomly alter blur orientations $[x;y]$ to $[-x;y]$, $[x;-y]$, or $[-x;-y]$ to augment blurred images with different orientations, achieving a similar performance improvement as well.
@@
In ID-Blau, we adopt the diffusion process in [12] to generate blurred versions of sharp images with specified blur conditions. However, without the diffusion process, we can still utilize the same reblurring model to learn the mapping from a sharp image plus a computed blur condition map to the corresponding blurred image with the typical L1 loss function. To analyze the impact of ID-Blau with or without the diffusion process on deblurring performance, in Table 3, we compare Baseline (the same as in Figure 8), “w/ ID-Blau," (ID-Blau without diffusion), and “w/ ID-Blau” (our model). The results show that “w/ ID-Blau” and “w/ ID-Blau” improve the Baseline by $0.57$dB and $0.80$dB, respectively, and the diffusion model achieves $0.23$dB additional performance gain.
@@
Table 3: Effectiveness of using a diffusion model in ID-Blau on MIMO-UNet’s performance in PSNR (dB), where “ID-Blau” denotes the proposed ID-Blau without its diffusion process.
Model Baseline w/ ID-Blau w/ ID-Blau
@@
Table 4: Comparison between ID-Blau and BGAN regarding the DBGAN’s performance in PSNR (dB), where ID-Blau denotes the proposed ID-Blau without its diffusion process.
Model Baseline w/ BGAN w/ ID-Blau w/ ID-Blau
PSNR 30.43 30.92 31.57 31.67
@@
Besides, we compare the performance gains of ID-Blau and that of BGAN [42], a representative reblurring GAN for augmenting blurred images to improve the DBGAN deblurring model. To fairly compare ID-Blau with BGAN, we utilize ID-Blau to improve DBGAN’s deblurring performance. In Table 4, “Baseline” denotes DBGAN’s performance trained on GoPro, and “w/ BGAN” denotes DBGAN’s performance with BGAN-augmented data. “w/ ID-Blau” and “w/ ID-Blau” denote DBGAN’s performance using ID-Blau-augmented data without and with the diffusion process, respectively. Note that GAN-based methods tend to generate unpredictable and uncontrollable blurred images, whereas ID-Blau can generate controllable results based on specified blur conditions. As a result, “w/ ID-Blau” and “w/ ID-Blau” improve the PSNR performance of DBGAN by $1.14$dB and $1.24$dB. Thanks to the proposed continuous blur condition map with ID-Blau, the simplified “w/ ID-Blau” without the diffusion process can still offer an additional gain of $0.47$dB over that with “w/ BGAN.” This verifies the effectiveness of ID-Blau compared to the GAN-based reblurring model BGAN.
@@ 5 Conclusion
6 Acknowledgments
This work was supported in part by the National Science and Technology Council (NSTC) under grants 112-2221-E-A49-090-MY3, 111-2628-E-A49-025-MY3, 112-2634-F-002-005, 112-2221-E-004-005, and 113-2923-E-A49-003-MY2. This work was funded in part by Qualcomm through a Taiwan University Research Collaboration Project.